OpenStudy (abbycross167):

Can someone please help me with this question? |- 3 ½ | - | 12 ¼

OpenStudy (abbycross167):

here is my answer~ |- 3 ½ | - | 12 ¼ | ~ Keep the -13 1/2 , change the subtraction sign to an addition sign, and change the 12 ¼ |-3 ½| + |12 ¼ | |8 ¾| 8 ¾ is the answer

OpenStudy (abbycross167):

My teacher said it was wrong....

OpenStudy (abbycross167):

@hartnn could you try to help me?

OpenStudy (skullpatrol):

What is the definition of | a |?

OpenStudy (abbycross167):

I'm not too sure

OpenStudy (skullpatrol):

it is always positive

OpenStudy (skullpatrol):

| 0 | = 0 because 0 is neither positive nor negative

OpenStudy (skullpatrol):

The opposite of a number a is written as -a. The positive number of any pair of opposite nonzero real numbers is called the absolute value of each number in the pair. The absolute value of a number a is written as | a |.

OpenStudy (abbycross167):

So -3 1/2 = 3 1/2?

OpenStudy (skullpatrol):

Yes

OpenStudy (skullpatrol):

If a is positive, | a | = a. If a is negative, | a | = –a. If a = 0, | a | = 0.

OpenStudy (abbycross167):

okay so |-3 1/2| - |12 1/4| = |3 1/2| - |-12 1/4| ? @skullpatrol

OpenStudy (skullpatrol):

Nope, do not change the sign of 12 1/4 because it is already positive. Just change the sign of -3 1/2 to 3 1/2 as you did.

OpenStudy (abbycross167):

oh ok then what?

OpenStudy (skullpatrol):

Perform the subtraction.

OpenStudy (skullpatrol):

| -3 1/2 | = 3 1/2 | 12 1/2 | = 12 1/4

OpenStudy (abbycross167):

3 1/2 - 12 1/4 = -8 3/4

OpenStudy (abbycross167):

@skullpatrol

OpenStudy (skullpatrol):

You got it! :D

OpenStudy (abbycross167):

awesome! is that the answer?

OpenStudy (skullpatrol):

Yup

OpenStudy (abbycross167):

great! thank you!!

OpenStudy (skullpatrol):

Thanks for trying to learn :-)